A solution contains 0.675 g of Ca2 in enough water to make a 1925 mL solution. What is the milliequivalents of Ca2 per liter (mEq/L) for this solution?
1
Expert's answer
2018-03-10T03:52:07-0500
n=m/M*2=0.675g/40*2=0.03375mol
c=n/V*1000=0.03375/1.925l*1000=17.5mEq/L